OpenStudy (anonymous):

Find the area bounded by f(x)= -x^2+x+3 and g(x)=2x + 1

OpenStudy (anonymous):

do you know which one is on top

OpenStudy (anonymous):

i believe f(x)= -x^2 + x+3

OpenStudy (mathteacher1729):

Do you know where the functions intersect? (are you allowed to use a graphing program to view the functions?)

OpenStudy (anonymous):

No. I am not sure where they intersect

OpenStudy (mathteacher1729):

To find where they intersect, set the functions equal to one another: In other words, solve -x^2+x+3 =2x + 1 (it's a quadratic which factors nicely). :)

OpenStudy (anonymous):

f(x)=g(x) you will find where they intersect

OpenStudy (across):

\[f(x)=-x^2+x+3,\]\[g(x)=2x+1.\] \[f(x)=g(x),\]\[-x^2+x+3=2x+1,\]\[-x^2-x+2=0\implies(-x+1)(x+2)=0.\] \[\int_{-2}^{1}\int_{2x+1}^{-x^2+x+3}dydx.\]

OpenStudy (anonymous):

is it 4/3 then across?

OpenStudy (anonymous):

Mathteacher1729: if i factor it i get 2 and 1.

OpenStudy (across):

I got 9/2.

OpenStudy (mathteacher1729):

Wow, across, that is an interesting solution. :D I'm willing to wager that treyhud's class has not yet covered double integrals though... Treyhud, what was the factorization you obtained? (the roots you gave are not quite accurate).

OpenStudy (anonymous):

whew again double integrals